# Josephson diodes induced by the loop current states

### v1

@@ Abstract
Introduction.— Superconducting diode effect (SDE) refers to the phenomenon where critical supercurrents flowing in opposite directions have different magnitudes. This non-reciprocal transport effect in the superconducting system can play a similar role as the diodes formed by the p-n junctions in the semiconductor industry and has potential applications in the development of dissipationless electronics [1, 2].
@@
SDE has been observed experimentally in numerous superconducting systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Such diode effect is also realized in the various Josephson junction (JJ) systems [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], which is named the Josephson diode effect (JDE). It provides an extra knob to realize and tune the diode effect by designing the proper barrier layer of the JJs.
@@
On the side of theoretical development, JDE was first proposed based on the electron and hole-doped superconductors (SC) in analog to the semiconducting p-n junctions [25]. From the symmetry point of view, both SDE and JDE require the system to break both inversion and time-reversal symmetry. These special JJs with both symmetry broken were intensively studied [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. One class of such JJs is called Josephson junction, which provides a possible mechanism to realize the JDE. More recently, as stimulated by the experimental observations, numerous theoretical proposals [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65] have been suggested for these effects. Moreover, the diode effect for the spin transport in the JJs is also proposed [66, 67].
@@
While most of the realizations of the SDE/JDE require a finite external or internal magnetic field to break the time-reversal symmetry, recent observations of the field-free diode effect in the JJs [14] and SC bulk [7] significantly expand the potential scope for this phenomenon but at the same time call for new theoretical mechanisms to realize it. Besides the systems with the magnetic field, the loop current states are distinct phases of matter that break the time-reversal symmetry, which are extensively studied in different systems, ranging from the Haldane model [68] on the honeycomb lattice to the various forms of flux phases in the square lattice [69, 70, 71, 72, 73, 74, 75] and triangular lattice [76, 77, 78, 79]. The recent discovery of the exotic charge density wave in nonmagnetic AV3Sb5 (A = K, Rb, Cs) [80, 81, 82, 83, 84, 85, 86, 87, 88] provides another possible material realization of the loop current states [89, 90, 91, 92, 93, 94, 95, 96, 97]. In this work, we propose a theoretical model to realize the JDE using the loop current states as the tunneling barrier of the JJ, which serves as the source of the time-reversal symmetry breaking. We construct the loop current states in the honeycomb lattice with monolayer and bilayer systems. For the monolayer case, JDE can be realized either by introducing a potential difference for the two sublattices or by modifying the current directions of the Haldane model that breaks the inversion symmetry directly. For the AB-staked bilayer case, JDE can be induced by either stacking the two layers with opposite loop currents or applying an out-of-plane electric field. For the bilayer system, we further analyze the effect of the interlayer coupling and displacement field on the diode efficiency.
@@
Tight-binding Models.— As the basic building block of the tunneling barrier of the JJ, we start with the Haldane model (HM) [68] on a honeycomb lattice containing two different types of atoms denoted A and B, as shown by the solid line in Fig.1(a). The tight-binding Hamiltonian can be written as
@@
where labels the lattice sites and is the electron spin index. The operators represent the annihilation (creation) operators of fermions located at site with spin . Moreover, is the nearest-neighbor (NN) hopping and denotes the amplitude of the next-nearest-neighbor (NNN) hopping. The complex NNN hopping breaks time-reversal symmetry by forming a loop current state, with , where and are the unit vectors along the two bonds that constitute the NNN bonds from site to site . The signs of the NNN hopping phase are schematically shown in Fig. 1(b), which is positive along the arrow direction. From the sign pattern, we can see that the bonds along the same direction, but consisting of different types of atoms, have hoppings with opposite phases, which preserves the inversion symmetry. Since the inversion symmetry has to be broken to achieve the JDE, we further consider the modified Haldane model (MHM) [98], where the sign of the hopping phase of the bonds consisting of atom B are reversed, as illustrated in Fig.1(c). Therefore, in the MHM, both time-reversal and inversion symmetry are broken, which meets the basic symmetry requirement to realize the JDE.
@@
Besides the monolayer system, we also consider the bilayer HM, where the two layers of the honeycomb lattice are Bernal stacked, with one layer’s A atoms stacked with the other layer’s B atoms as shown in Fig.1(a). The Hamiltonian for the bilayer system can thus be written as
@@
where is the layer index, represents the Hamiltonian for each layer, and corresponds to the coupling between the two layers. Here, we assume that only the atoms stacked on top of each other between the two layers (A1-B2) contribute to this interlayer coupling . The introduction of the extra layer brings more freedom to manipulate the symmetry of the system. Here, we consider the loop current pattern of each layer forms the original HM, which is demonstrated to be more stable in a mean-field calculation [99]. Then, the remaining degree of freedom is the relative orientation of the loop current of the two layers. When the two layers have the loop current with the same orientation as depicted in Fig. 1(d), the inversion symmetry remains preserved and the bilayer system has similar properties as the monolayer HM. In order to break the inversion symmetry in this state, an electric field perpendicular to the layers can be introduced. On the other hand, when the orientations of the loop current are opposite for each layer as shown in Fig. 1(e), the inversion symmetry is broken by having an odd parity with respect to two layers. This is the magnetoelectric state theoretically proposed in the bilayer honeycomb lattice with NNN repulsion [99].
@@
Construction of the Josephson junction.— We next construct the planar Josephson junction using the monolayer and bilayer systems as the tunneling barrier. As depicted in Fig.2, this can be achieved by placing two s-wave superconductors (blue) on a two-dimensional material with loop current order (red). Due to the proximity effect, the two-dimensional material underneath the superconductors becomes superconducting. The region without superconductors in between remains in the normal state, thus forming a Josephson junction. Since the translation symmetry is preserved in the y-direction, we adopt the periodic boundary condition in this direction. In the y-direction, we consider unit cells , while in the x-direction, we have 20 layers for the left superconducting region, 2 layers for the middle normal region, and 20 layers for the right superconducting region.
@@
where represents the Hamiltonian of monolayer or bilayer system, as derived from either Eq.1 or Eq.2. and are the Hamiltonians of the superconducting regions. represents the Hamiltonian of the non-superconducting region. and represent the coupling terms between them. The superconducting phase bias for the JJ is assigned to the left superconducting region as so that the phase in the right superconducting region is set to . The amplitude of the order parameter is taken as for all calculations in this paper. represents the chemical potential of the entire material, which is set to -0.4t in the bilayer model, and set to 0 in the monolayer system.
@@
where represents the th eigenvalue of the total Hamiltonian and denotes the Fermi-Dirac distribution function, with the temperature set to be through the whole paper. In Eq. 4, while considering the contribution of the superconducting current, all states are included.
@@
The current-phase relationship.— We construct the JJs with both the monolayer and bilayer systems as the tunneling barrier. Depending on the orientation of the junction, we name the JJ as a zigzag (armchair) JJ, if the edge of the junction along the current flow direction has a zigzag (armchair) shape. For the monolayer case, when the tunneling barrier is constructed with the HM, the current-phase relation (CPR) shows no diode effect, where the critical current for the positive and negative directions have the same magnitude [101]. This is expected since the loop current pattern in the HM preserves the inversion symmetry, which precludes the emergence of the JDE. One way to break this inversion symmetry is to introduce a staggered lattice potential within the unit cell of HM, which can be described by an extra term . The CPR for both types of the JJ is shown in Fig. 3(a, b). Indeed, we can see that the zigzag JJ shows the diode effect where the magnitude of the critical current for both directions differs from each other, i.e., as shown in Fig. 3(a). Moreover, a finite Josephson current is achieved for the vanishing phase bias , which is directly related to the breaking of both inversion and time-reversal symmetry in this system. However, the CPR for the armchair JJ does not show any nonreciprocal effect. We can understand the different behaviors of the two types of JJs from the symmetry argument. From the symmetry point of view, if the system is invariant under a certain operation that also changes the sign of the current operator, the state of the system with a positive current is always connected to another state with its current reversed by this operation. This means the critical current must have the same magnitude for opposite directions, which causes the absence of the JDE. Since both time-reversal and inversion change the sign of the current operator, the JDE requires both symmetries to be broken. However, even in a system with both symmetries broken, as long as there exists a symmetry that changes the sign of the current operator, the JDE is still strictly forbidden. For the armchair JJ with staggered potential, this symmetry is , where is the mirror reflection about the plane perpendicular to the armchair edge, is the 2-fold rotation about the axis perpendicular to the JJ and is the time-reversal operator. Since all three operations change the current direction of the armchair JJ, this combined symmetry operation does change the sign of the current operator in the armchair JJ and thus forbids the JDE.
On the other hand, if the tunneling barrier is constructed with the MHM, where the inversion symmetry is explicitly broken by the loop current pattern, we also expect the JDE to show up. As shown in Fig. 3(c, d), we again find that the JDE only shows up for the zigzag JJ. This time, it is the symmetry that prevents the JDE from taking place in the armchair JJ.
@@
We next consider the JJs constructed with the bilayer HM. Based on the analysis of the monolayer case, we can see that breaking the inversion symmetry becomes the key factor in realizing the JDE in the system since the time-reversal symmetry is already broken. We first consider the case where the orientation of the loop current for the two layers is the same, which still preserves the inversion symmetry and thus prevents the emergence of the JDE. In this system, the inversion symmetry can be broken by applying an external electric field perpendicular to the layers, resulting in a potential difference between the two layers. In our calculation, this effect is described by a layer-dependent potential term which is added to the tunneling barrier part of the Hamiltonian in Eq. 3. In this case, the JDE again shows up only for the zigzag JJ. The CPR for the zigzag JJ with is shown in Fig. 4(a), which clearly shows the JDE induced by the external electric field.
@@
Besides the case where the inversion symmetry is broken by the external field, we next consider the bilayer magnetoelectric state where the orientations of the loop current of the two layers are opposite to each other so that the inversion symmetry is broken intrinsically. In this case, the JDE can be realized without the help of an external field as shown in Fig. 4(b) for the CPR of the zigzag JJ. The CPR for the armchair JJ again does not show any diode effect. A thorough symmetry analysis for the absence of the JDE in the armchair JJ is provided in the supplemental material [101].
@@
Tuning the diode efficiency.— To investigate the factors influencing the diode effect, we use the diode efficiency factor to quantify the extent of the diode effect. It is defined as:
@@
where is the critical current along the positive (negative) direction, and reflects the strength of the diode effect. For the bilayer HMs, the inversion symmetry is broken by applying the interlayer voltage U for the case with the same loop current direction and the interlayer hopping for the case with opposite directions of the loop current breaks the independent rotation symmetry, both of which result in the JDE. Therefore, it is natural to investigate the effects of U and on the diode efficiency.

### v3

@@ I Introduction
Superconducting diode effect (SDE) refers to the phenomenon where critical supercurrents flowing in opposite directions have different magnitudes. This non-reciprocal transport effect in the superconducting system can play a similar role as the diodes formed by the p-n junctions in the semiconductor industry and has potential applications in the development of dissipationless electronicsΒ [1, 2].
@@
SDE has been observed experimentally in numerous superconducting systemsΒ [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Such diode effect is also realized in the various Josephson junction (JJ) systemsΒ [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], which is named the Josephson diode effect (JDE). It provides an extra knob to realize and tune the diode effect by designing the proper barrier layer of the JJs.
@@
On the side of theoretical development, JDE was first proposed based on the electron and hole-doped superconductors (SC) in analog to the semiconducting p-n junctionsΒ [25]. From the symmetry point of view, both SDE and JDE require the system to break both inversion and time-reversal symmetry. These special JJs with both symmetry broken were intensively studiedΒ [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. One class of such JJs is called Josephson junction, which provides a possible mechanism to realize the JDE. More recently, as stimulated by the experimental observations, numerous theoretical proposalsΒ [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70] have been suggested for these effects. Moreover, the diode effect for the spin transport in the JJs is also proposedΒ [71, 72].
@@
While most of the realizations of the SDE/JDE require a finite external magnetic field or internal net magnetic moment to break the time-reversal symmetry, recent observations of the field-free diode effect in the JJsΒ [14] and SC bulkΒ [7] significantly expand the potential scope for this phenomenon but at the same time call for new theoretical mechanisms to realize it. Besides the systems with the magnetic field, the loop current states are distinct phases of matter that break the time-reversal symmetry, which are extensively studied in different systems, ranging from the Haldane modelΒ [73] on the honeycomb lattice to the various forms of flux phases in the square latticeΒ [74, 75, 76, 77, 78, 79, 80] and triangular latticeΒ [81, 82, 83, 84]. The recent discovery of the exotic charge density wave in nonmagnetic AV3Sb5 (A = K, Rb, Cs)Β [85, 86, 87, 88, 89, 90, 91, 92, 93] provides another possible material realization of the loop current statesΒ [94, 95, 96, 97, 98, 99, 100, 101, 102]. In this work, we propose a theoretical model to realize the JDE using the loop current states as the tunneling barrier of the JJ, which serves as the source of the time-reversal symmetry breaking. We construct the loop current states in the honeycomb lattice with monolayer and bilayer systems. For the monolayer case, JDE can be realized either by introducing a potential difference for the two sublattices or by modifying the current directions of the Haldane model that breaks the inversion symmetry directly. For the AB-staked bilayer case, JDE can be induced by either stacking the two layers with opposite loop currents or applying an out-of-plane electric field. For the bilayer system, we further analyze the effect of the interlayer coupling and displacement field on the diode efficiency.
This work is organized as follows. In Sec.II, we introduce the tight-binding models that host the loop current order. In Sec.III, we describe the construction of the Josephson junction within the tight-binding model. The main results, including the current-phase relation of the junction, are presented in Sec.IV, followed by an analysis of the diode efficiency in Sec.V. In Sec.VI, we provide a symmetry analysis of the Josephson diode effect (JDE), and in Sec.VII, we examine its robustness against disorder. Finally, we conclude our discussion in Sec.Β VIII.
II Tight-binding Models
@@
As the basic building block of the tunneling barrier of the JJ, we start with the Haldane model (HM)Β [73] on a honeycomb lattice containing two different types of atoms denoted A and B, as shown by the solid line in Fig.1(a). The tight-binding Hamiltonian can be written as
@@
where labels the lattice sites and is the electron spin index. The operators represent the annihilation (creation) operators of fermions located at site with spin . Moreover, is the nearest-neighbor (NN) hopping and denotes the amplitude of the next-nearest-neighbor (NNN) hopping. The complex NNN hopping breaks time-reversal symmetry by forming a loop current state, with , where and are the unit vectors along the two bonds that constitute the NNN bonds from site to site . The signs of the NNN hopping phase are schematically shown in Fig.Β 1(b), which is positive along the arrow direction. From the sign pattern, we can see that the bonds along the same direction, but consisting of different types of atoms, have hoppings with opposite phases, which preserves the inversion symmetry. Since the inversion symmetry has to be broken to achieve the JDE, we further consider the modified Haldane model (MHM)Β [103], where the sign of the hopping phase of the bonds consisting of atom B are reversed, as illustrated in Fig.1(c). Therefore, in the MHM, both time-reversal and inversion symmetry are broken, which meets the basic symmetry requirement to realize the JDE.
@@
Besides the monolayer system, we also consider the bilayer HMs, where the two layers of the honeycomb lattice are Bernal stacked, with one layerβs A atoms stacked with the other layerβs B atoms as shown in Fig.1(a). The Hamiltonian for the bilayer system can thus be written as
@@
where is the layer index, represents the Hamiltonian for each layer, and corresponds to the coupling between the two layers. Here, we assume that only the atoms stacked on top of each other between the two layers (A1-B2) contribute to this interlayer coupling . The introduction of the extra layer brings more freedom to manipulate the symmetry of the system. Here, we consider the loop current pattern of each layer forms the original HM, which is demonstrated to be more stable in a mean-field calculationΒ [104]. Then, the remaining degree of freedom is the relative orientation of the loop current of the two layers. When the two layers have the loop current with the same orientation as depicted in Fig.Β 1(d), the inversion symmetry remains preserved and the bilayer system has similar properties as the monolayer HM. In order to break the inversion symmetry in this state, an electric field perpendicular to the layers can be introduced. On the other hand, when the orientations of the loop current are opposite for each layer as shown in Fig.Β 1(e), the inversion symmetry is broken by having an odd parity with respect to two layers. This is the magnetoelectric state theoretically proposed in the bilayer honeycomb lattice with NNN repulsionΒ [104].
III Construction of the Josephson junction
@@
We next construct the planar Josephson junction using the monolayer and bilayer systems as the tunneling barrier. As depicted in Fig.2, this can be achieved by placing two s-wave superconductors (blue) on a two-dimensional material with loop current order (red). Due to the proximity effect, the two-dimensional material underneath the superconductors becomes superconducting. The region without superconductors in between remains in the normal state, thus forming a Josephson junction. Since the translation symmetry is preserved in the y-direction, we adopt the periodic boundary condition in this direction. In the y-direction, we consider unit cells , while in the x-direction, we have 20 layers for the left superconducting region, 2 layers for the middle normal region, and 20 layers for the right superconducting region.
@@
where represents the Hamiltonian of monolayer or bilayer system, as derived from either Eq.S2 or Eq.S3. and are the Hamiltonians of the superconducting regions. represents the Hamiltonian of the non-superconducting region. and represent the spin-conserving hoppings between the superconducting and normal sections which include both the NN and NNN hoppings and take the same values and patterns as those in the normal region. The superconducting phase bias for the JJ is assigned to the left superconducting region as so that the phase in the right superconducting region is set to . The amplitude of the order parameter is taken as for all calculations in this paper. represents the chemical potential of the entire material, which is set to -0.4t in the bilayer model, and set to 0 in the monolayer system. More detailed setups for the tight-binding model of the JJs are provided in the Supplemental MaterialΒ [105].
@@
where represents the th eigenvalue of the total Hamiltonian and denotes the Fermi-Dirac distribution function, with the temperature set to be through the whole paper. In Eq.Β 4, while considering the contribution of the superconducting current, all states are included.
IV The current-phase relationship
@@
We construct the JJs with both the monolayer and bilayer systems as the tunneling barrier. Depending on the orientation of the junction, we name the JJ as a zigzag (armchair) JJ, if the edge of the junction along the current flow direction has a zigzag (armchair) shape. For the monolayer case, when the tunneling barrier is constructed with the HM, the current-phase relation (CPR) shows no diode effect, where the critical current for the positive and negative directions have the same magnitudeΒ [105]. This is expected since the loop current pattern in the HM preserves the inversion symmetry, which precludes the emergence of the JDE. One way to break this inversion symmetry is to introduce a staggered lattice potential within the unit cell of HM, which can be described by an extra term
This term is added to Eq.Β S2 which enters the Hamiltonian of the JJ as in Eq.Β S1. The staggered potential can be realized by placing the material on a substrate with a honeycomb structure, such as Hexagonal boron nitride (h-BN), where the two sublattices consist of different atoms. The CPR for both types of the JJ is shown in Fig.Β 3(a, b). Indeed, we can see that the zigzag JJ shows the diode effect where the magnitude of the critical current for both directions differs from each other, i.e., as shown in Fig.Β 3(a). Moreover, a finite Josephson current is achieved for the vanishing phase bias , which is directly related to the breaking of both inversion and time-reversal symmetry in this system. However, the CPR for the armchair JJ does not show any nonreciprocal effect. We can understand the different behaviors of the two types of JJs from the symmetry argument. From the symmetry point of view, if the system is invariant under a certain operation that also changes the sign of the current operator, the state of the system with a positive current is always connected to another state with its current reversed by this operation. This means the critical current must have the same magnitude for opposite directions, which causes the absence of the JDE. Since both time-reversal and inversion change the sign of the current operator, the JDE requires both symmetries to be broken. However, even in a system with both symmetries broken, as long as there exists a symmetry that changes the sign of the current operator, the JDE is still strictly forbidden. The detailed symmetry analysis will be presented in Sec.Β VI. On the other hand, if the tunneling barrier is constructed with the MHM, where the inversion symmetry is explicitly broken by the loop current pattern, we also expect the JDE to show up. As shown in Fig.Β 3(c, d), we again find that the JDE appears only for the zigzag JJ due to similar symmetry reasons.
@@
We next consider the JJs constructed with the bilayer HMs. Based on the analysis of the monolayer case, we can see that breaking the inversion symmetry becomes the key factor in realizing the JDE in the system since the time-reversal symmetry is already broken. We first consider the case where the orientation of the loop current for the two layers is the same, which still preserves the inversion symmetry and thus prevents the emergence of the JDE. In this system, the inversion symmetry can be broken by applying an external electric field perpendicular to the layers, resulting in a potential difference between the two layers. In our calculation, this effect is described by a layer-dependent potential term
which is added to the tunneling barrier part of the Hamiltonian in Eq.Β S1. In this case, the JDE again shows up only for the zigzag JJ. The CPR for the zigzag JJ with is shown in Fig.Β 4(a), which clearly shows the JDE induced by the external electric field. The CPR for the armchair JJ is provided in the Supplemental MaterialΒ [105].
@@
Besides the case where the inversion symmetry is broken by the external field, we next consider the bilayer magnetoelectric state where the orientations of the loop current of the two layers are opposite to each other so that the inversion symmetry is broken intrinsically. In this case, the JDE can be realized without the help of an external field as shown in Fig.Β 4(b) for the CPR of the zigzag JJ. The CPR for the armchair JJ again does not show any diode effectΒ [105]. A thorough symmetry analysis for the absence of the JDE in the armchair JJ is provided in Sec.Β VI.
V Tuning the diode efficiency
@@
To investigate the factors influencing the diode effect, we use the diode efficiency factor to quantify the strength of the diode effect. It is defined as:
@@
where is the critical current along the positive (negative) direction. For the bilayer HMs, the inversion symmetry is broken by applying the interlayer voltage U for the case with the same loop current direction and the interlayer hopping for the case with opposite directions of the loop current breaks the independent rotation symmetry, both of which result in the JDE. Therefore, it is natural to investigate the effects of U and on the diode efficiency.
@@
The diode efficiency is sensitive to a range of parameters, including the interlayer voltage , the interlayer coupling , the chemical potential , the pairing order parameter , the flux phase , the size of the system, and so on. Only in the parameter region where the parameters responsible for the symmetry breaking are small, the dependence is smooth and can be understood in the spirit of the perturbation theory. For the bilayer case, we also explore the effect the chemical potential and the pairing amplitude to , in the wide range of and with the results shown in the Supplemental materialΒ [105]. The strong sensitivity of to the parameters requires precise experimental control when implementing the effect in realistic systems.
VI Symmetry analysis
In this section, we demonstrate the reason why the JDE realized in the zigzag JJs disappears in the armchair JJs. From the symmetry perspective, if the system remains invariant under a certain operation that simultaneously changes the sign of the current operator, it implies that the critical current must have the same magnitude in opposite directions, resulting in the absence of the JDE. We find that such symmetry operations can be constructed by the following three operations: the mirror reflection (), the time-reversal (), and the rotation about the axis perpendicular to the plane. Both and operations can change the current direction. represents the mirror reflection about the plane perpendicular to the -axis, and the definition of the directions is schematically shown in Fig.Β 5, where corresponds to the zigzag direction and represents the armchair direction. For the zigzag JJ, changes the current direction while does not. Conversely, for the armchair JJ, changes the current direction, but does not.
For JJs constructed by different models with different junction directions, the possible symmetry operations that prevent the emergence of the JDE are summarized in Table.Β 1. For example, for the armchair JJ constructed by the monolayer HM model with staggered lattice potential, operations , and all change the direction of the current. Therefore, the combined operation changes the direction of the current while keeping the system invariant. Hence, the JDE is absent in the armchair JJ constructed by the HM model with staggered lattice potential. For zigzag JJ constructed by the same model, we can not find any symmetry operations that also change the direction of the current which is represented by the ββ sign in the table.
| monolayer I | operations | JDE | control strategy |
| armchair HM | No | / | |
| armchair MHM | No | ||
| zigzag HM | No | ||
| zigzag MHM | Yes | ||
| monolayer II | operations | JDE | control strategy |
| armchair HM | No | place the system on a substrate of h-BN | |
| armchair MHM | No | ||
| zigzag HM | Yes | ||
| zigzag MHM | Yes | ||
| bilayer I | operations | JDE | control strategy |
| armchair HMs | No | apply an external electric field perpendicular to the barrier | |
| zigzag HMs | Yes | ||
| bilayer II | operations | JDE | control strategy |
| armchair HMs | No | / | |
| zigzag HMs | Yes |
VII The impact of disorder
To investigate the impact of impurities on the diode effect, we study the JDE with a random on-site potential describing the disorder effect introduced by the impurities. We use the random potentials obeying the uniform distribution ranging from to and the results are averaged over 100 disorder configurations. Since the translation invariance along the -direction is broken by the random impurity potential, we perform the calculation in real space, which significantly increases the computational complexity. We therefore reduce the model size. The size in the -direction remains unchanged, while in the -direction, it is reduced from 100 unit cells (4100 atoms) to 10 unit cells (410 atoms). The results for the zigzag JJ constructed by the single-layer MHM are shown in Fig.Β 6, where the averaged CPR and diode efficiency obtained from the disordered model with various disorder strengths are compared to those of the impurity-free model.
We perform the calculation for two values of the superconducting order parameter with = and . The results clearly show that the CPRs for the disordered JJs still show the diode effect and the diode efficiency persists with the increase of the disorder strength. Moreover, for the relatively strong pairing strength , the critical currents are less affected by the disorder strength as evidenced by the close resemblance of the CPRs across different disorder strengths, leading to a stable diode efficiency for . It starts to decrease slightly for above . On the other hand, for the relatively weak pairing strength , the supercurrents become weaker and more susceptible to disorder, which manifests as a rapid decay of the critical supercurrents in both directions as shown in Fig.Β 6(b). However, since the breaking of time-reversal and inversion symmetries is governed by the loop currents rather than , the diode effect become more pronounced in this case and starts to decrease for above . This rules out the possibility that the βcriticalβ disorder strength where starts to decrease is restricted by , thereby indicating that the diode effect is robust against the disorder.
The robustness of the JDE against disorder can also be understood as follows. The diode efficiency defined in Eq.Β 7 quantifies the relative difference in critical supercurrents flowing in opposite directions. Its physical origin lies in the breaking of certain symmetries, such as time-reversal and inversion symmetries, that otherwise relate the current in opposite directions. In this sense, serves as a measure of symmetry breaking. Consequently, it is reasonable to expect that remains robust when the disorder is not too strong, as it cannot restore the broken symmetries responsible for the diode effect.
VIII Discussion